Ques 1. Find the y-intercept of the graph of 6-3y+x=0
Ques 2. What are the x-intercepts of y=20+3(2x+1)
Ques 3. Find the equation of the line that is parallel fo the graph of y=3x-2 and has the y-intercept of (0,3).
Ques 4. Solve by completing the square. Lowest possible terms: x^2-8x+2=0
Ques 5. Solve by completing the square. Lowest possible terms: 6x-x^2=3
Ques 6. Solve for x - Lowest possible terms: 2x^2=162
Ques 7. Solve for x - lowest possible terms: (2x+3)^2=4
Ques 8. Solve for x - lowest possible terms: 5x^2-2=3x

Answer by sehar(13)

solution 1:
to find y-intercept, put x=0
now,
6-3y+0=0
6-3y=0
6=3y
or,
3y=6
y=6/3
y=2
so, y-intercept is 2.
solution 2:
here,
y= 20+3(2x+1)
y=20+6x+3
to find x-intercept,put y=0
0=23+6x
6x=-23
x= -23/6
solution 3:
the given line is y= 3x-2
let the required line be y=mx+c, where m is slope
slope of given line is (-coeficient of x/coefficient of y) i.e. -3
parallel lines hv equal slopes always so put m=-3 in y=mx+c
y=-3x+c.....(i)
now as (0,3) pass through the reuired line ,it must satisfy (i)
3=0+c
so, c=3
put it in (i)
y=-3x+3
which is the required equation